# Street Gaussians without 3D Object Tracker

### v2

@@ 1 Introduction
Modeling dynamic 3D street scenes underpins modern autonomous driving by enabling realistic, controllable simulations for tasks such as perception [47, 11, 83, 39], prediction [31, 43, 23], and motion planning [13, 14, 16]. With the rise of end-to-end autonomous systems that require real-time sensor feedback [27, 29, 32], high-quality scene reconstructions have become essential for closed-loop evaluations [85, 42], particularly to simulate critical corner cases safely and cost-effectively.
Despite extensive efforts in achieving photo-realistic reconstruction of small-scale scenes, the unique challenges posed by the large-scale and highly dynamic nature of driving scenarios complicate effective 3D scene modeling. To address these challenges, most existing methods [35, 74, 90, 56] rely on ground truth vehicle poses to differentiate between static background and moving vehicles. Typically, vehicles are reconstructed in canonical space and subsequently positioned based on known poses during rendering. However, collecting ground-truth poses is labor-intensive, limiting the applicability of these methods to scenes beyond the existing datasets.
@@ 2 Related Works
3D Gaussian Splatting (3DGS) [34] has advanced scene reconstruction by enabling high-quality, real-time rendering with 3D Gaussians and efficient splat-based rasterization, reducing computation and parameters compared to NeRF-based methods [52, 1, 2, 3, 54] and other representations [15, 24, 65, 25, 28, 41]. While originally designed for static scenes, 3DGS has been adapted for dynamic scenes [48, 81, 82, 70, 64, 44]. Dynamic3DGS [48] directly stores information for each 3D Gaussian at every timestamp, and Yang et al. [81] approximate the spatiotemporal 4D volume by optimizing 4D Gaussian primitives. Deformable-3DGS [82] and GauFRe [44] employ deformation fields to model motions, while 4DGS [70] introduces HexPlane representation [7] to store spatial-temporal features efficiently. Gaussian Marbles [64] leverages isotropic “marbles” and a divide-and-conquer trajectory learning algorithm. Despite these advances, challenges persist in high-speed driving scenarios with rapid object motion. To tackle this challenge, we leverage a robust object tracking module to guide the motion learning process.
@@ 2.2 Street Scene Reconstruction
Existing autonomous driving simulation engines [19, 61, 40] face high manual effort in creating virtual environments and a lack of realism in the generated data. The creation of high-fidelity simulations from driving logs is therefore essential for advancing closed-loop training and testing. Recent works [80, 76, 56, 67, 62, 74, 79] have continuously made improvements to NeRF [52] to model street scenes dynamically. Despite the progress, NeRF-based methods remain computationally expensive and require densely overlapping views. Building on the effective 3DGS approach [34] for scene reconstruction, several 3DGS-based methods have emerged. PVG [10] model dynamic scenarios by using periodic vibration-based temporal dynamics. Driving Gaussian [90] introduces incremental static Gaussians and composite dynamic Gaussian graph. Street Gaussians [78] equips Gaussians with semantic logits, and optimize dynamic parts using tracked poses from 3D tracker. AutoSplat [35] enforces geometric constraints in road and sky regions for multi-view consistency. Among these approaches, most of them requires ground truth object pose [10, 90, 35] or the 3D tracker [78]. However, manual annotation is laborious and the 3D tracker is lack in generalization ability, which limits their applications in diverse scenarios. In contrast, S3Gaussian [30] models dynamics in a self-supervised manner using dynamic Gaussians from 4DGS [70], despite struggling with dynamic object modeling due to the lack of explicit motion supervision. We address this challenge by introducing a robust object tracking strategy based on a 2D foundation model [73] and apply motion supervision from the predicted object trajectory.
@@
Multiple Object Tracking (MOT) aims at locating multiple objects in each frame, and establishing correspondences between them across frames within input videos [4, 87, 49, 33, 5]. The reliance of most MOT methods on object detection [4, 87, 49, 57, 18, 89] makes detection accuracy crucial for MOT performance. Recent advances in 3D object detection are promising [69, 51, 83, 11, 47, 39], but these models often generalize poorly due to limited and biased datasets [63, 86, 20]. Existing open-source autonomous driving datasets lack scale and show regional or environmental biases, such as vehicle density and weather conditions [66, 6, 21, 50]. The main reason is that building large-scale, well-annotated multimodal datasets is costly [6, 50, 59, 45]. In contrast, 2D image data is easier to capture and annotate, enabling extensive dataset collections [9, 84, 17, 38]. Visual foundation models trained on large-scale 2D data have shown strong generalization [73, 12, 58, 37, 55]. We thus opt to use 2D object tracker [73] to locate the dynamic objects.
@@ 3.3 Tracking Cars in 3D with Robust 2D Tracker
After obtaining the 3D partial objects in different time steps, the next step is to associate these partial point cloud into a unified and complete model. We initialize the 3D reconstruction of each object from the first frame which it appears. For simplicity, we assume an object is visible between frames 00 and $T$ and denote the object point cloud at time $t$ as $\mathcal{O}_{t}$. Our goal is to obtain a temporally consistent reconstruction $\mathcal{O}$ and the object pose in each time frame. This reconstruction is done frame-by-frame incrementally.
@@ 4.2 Comparisons with the State-of-the-art
We present qualitative comparison with S3Gaussian [30] and Street Gaussians [78] in Fig. 3. Results from S3Gaussian show that using 4DGS without explicit motion guidance results in weaker performance when handling moving vehicles (e.g. row 2, 3, 5). Street Gaussians suffers from tracking errors of 3D trackers, leading to erroneous reconstructions (e.g. row 1, 3) or entirely missed objects (e.g. rows 2, 4, and 5). In contrast, our approach performs robustly across diverse scenarios owing to our tracking strategy leveraging 2D foundation model and robust motion learning module.
@@ 4.3 Ablation Studies
Effect of motion learning strategy. In Tab. 3 (B), we use the reconstruction of the tracking module for initialization but remove the motion loss introduced in Eq. 7, resulting in a 1.04 dB drop in DPSNR. This result highlights the importance of our motion learning strategy.
@@
3D tracker / GT pose vs. Our object tracking module. In Tab. 3 (D), we use the 3D tracker [11, 57] to reconstruct dynamic objects and supervise motion learning, resulting in a 0.97 dB DPSNR decrease. Without the motion loss, DPSNR decreases further by 1.08 dB (Tab. 3 (E)), showing that our motion learning strategy can compensate for tracking errors, though errors still impact novel view synthesis, emphasizing the need for a more generalizable 2D tracker. In Tab. 3 (F), we conduct an experiment using the ground truth object trajectory to benchmark our 2D tracking approach. The relatively small performance gap indicates that our 2D tracking closely approximates the ground truth in our motion learning procedure, supporting robust performance across varied scenarios.
@@ 5 Conclusion
In this paper, we introduce a novel framework for robust dynamic 3D street scene reconstruction that eliminates the reliance on 3D object trackers. Addressing the generalization limitations of 3D trackers, we propose a robust object tracking strategy based on a 2D foundation model. Our framework also features a motion learning module within an implicit feature space to handle inevitable tracking errors by autonomously refining pose inaccuracies and recovering missed detections. Experiments on the Waymo-NOTR and KITTI datasets demonstrate its adaptability and superior performance.

### v3

@@ 1 Introduction
Modeling dynamic 3D street scenes underpins modern autonomous driving by enabling realistic, controllable simulations for tasks such as perception [56, 11, 96, 42, 107], prediction [33, 47, 25], and motion planning [13, 14, 16]. With the rise of end-to-end autonomous systems that require real-time sensor feedback [29, 31, 35], high-quality scene reconstructions have become essential for closed-loop evaluations [98, 46], particularly to simulate critical corner cases safely and cost-effectively.
Despite extensive efforts in achieving photo-realistic reconstruction of small-scale scenes, the unique challenges posed by the large-scale and highly dynamic nature of driving scenarios complicate effective 3D scene modeling. To address these challenges, most existing methods [38, 85, 108, 65, 22] rely on ground truth vehicle poses to differentiate between static background and moving vehicles. Typically, vehicles are reconstructed in canonical space and subsequently positioned based on known poses during rendering. However, collecting ground-truth poses is labor-intensive, limiting the applicability of these methods to scenes beyond the existing datasets.
@@ 2 Related Works
3D Gaussian Splatting (3DGS) [37, 44] has advanced scene reconstruction by enabling high-quality, real-time rendering with 3D Gaussians and efficient splat-based rasterization, reducing computation and parameters compared to NeRF-based methods [61, 1, 2, 3, 63] and other representations [15, 26, 75, 27, 30, 45, 104]. While originally designed for static scenes, 3DGS has been adapted for dynamic scenes [57, 94, 95, 81, 74, 48, 53]. Dynamic3DGS [57] directly stores information for each 3D Gaussian at every timestamp, and Yang et al. [94] approximate the spatiotemporal 4D volume by optimizing 4D Gaussian primitives. Deformable-3DGS [95] and GauFRe [48] employ deformation fields to model motions, while 4DGS [81] introduces HexPlane representation [7] to store spatial-temporal features efficiently. Gaussian Marbles [74] leverages isotropic “marbles” and a divide-and-conquer trajectory learning algorithm. Despite these advances, challenges persist in high-speed driving scenarios with rapid object motion. To tackle this challenge, we leverage a robust object tracking module to guide the motion learning process.
@@ 2.2 Street Scene Reconstruction
Existing autonomous driving simulation engines [20, 71, 52, 43] face high manual effort in creating virtual environments and a lack of realism in the generated data. The creation of high-fidelity simulations from driving logs is therefore essential for advancing closed-loop training and testing. Recent works [93, 87, 65, 78, 72, 85, 92, 77] have continuously made improvements to NeRF [61] to model street scenes dynamically. Despite the progress, NeRF-based methods remain computationally expensive and require densely overlapping views. Building on the effective 3DGS approach [37] for scene reconstruction, several 3DGS-based methods [10, 67, 108, 91] have emerged. PVG [10] model dynamic scenarios by using periodic vibration-based temporal dynamics. Driving Gaussian [108] introduces incremental static Gaussians and composite dynamic Gaussian graph. Street Gaussians [91] equips Gaussians with semantic logits, and optimize dynamic parts using tracked poses from 3D tracker. AutoSplat [38] enforces geometric constraints in road and sky regions for multi-view consistency. Among these approaches, most of them requires ground truth object pose [10, 108, 38] or the 3D tracker [91]. However, manual annotation is laborious and the 3D tracker is lack in generalization ability, which limits their applications in diverse scenarios. In contrast, S3Gaussian [32] models dynamics in a self-supervised manner using dynamic Gaussians from 4DGS [81], despite struggling with dynamic object modeling due to the lack of explicit motion supervision. We address this challenge by introducing a robust object tracking strategy based on a 2D foundation model [84] and apply motion supervision from the predicted object trajectory.
@@
Multiple Object Tracking (MOT) aims at locating multiple objects in each frame, and establishing correspondences between them across frames within input videos [4, 100, 58, 36, 5]. The reliance of most MOT methods on object detection [4, 100, 58, 66, 19, 106] makes detection accuracy crucial for MOT performance. Recent advances in 3D object detection are promising [80, 60, 96, 11, 56, 42, 18, 101, 102, 51, 103, 34, 54], but these models often generalize poorly due to limited and biased datasets [73, 99, 21, 55]. Existing open-source autonomous driving datasets lack scale and show regional or environmental biases, such as vehicle density and weather conditions [76, 6, 23, 59]. The main reason is that building large-scale, well-annotated multimodal datasets is costly [6, 59, 69, 49]. In contrast, 2D image data is easier to capture and annotate, enabling extensive dataset collections [9, 97, 17, 41]. Visual foundation models trained on large-scale 2D data have shown strong generalization [84, 12, 68, 40, 64]. We thus opt to use 2D object tracker [84] to locate the dynamic objects.
@@ 3.3 Tracking Cars in 3D with Robust 2D Tracker
After obtaining the 3D partial objects in different time steps, the next step is to associate these partial point cloud into a unified and complete model. We initialize the 3D reconstruction of each object from the first frame which it appears. For simplicity, we assume an object is visible between frames 0 and $T$ and denote the object point cloud at time $t$ as $\mathcal{O}_{t}$. Our goal is to obtain a temporally consistent reconstruction $\mathcal{O}$ and the object pose in each time frame. This reconstruction is done frame-by-frame incrementally.
@@ 4.2 Comparisons with the State-of-the-art
We present qualitative comparison with S3Gaussian [32] and Street Gaussians [91] in Fig. 3. Results from S3Gaussian show that using 4DGS without explicit motion guidance results in weaker performance when handling moving vehicles (e.g. row 2, 3, 5). Street Gaussians suffers from tracking errors of 3D trackers, leading to erroneous reconstructions (e.g. row 1, 3) or entirely missed objects (e.g. rows 2, 4, and 5). In contrast, our approach performs robustly across diverse scenarios owing to our tracking strategy leveraging 2D foundation model and robust motion learning module. Please refer to the supplementary material for additional video results and editing examples.
@@ 4.3 Ablation Studies
Effect of motion learning strategy. In Tab. 3 (B), we use the reconstruction of the tracking module for initialization but remove the motion loss introduced in Eq. 7, resulting in a 1.04 dB drop in DPSNR. This result highlights the importance of our motion learning strategy. A comparison with rigid-transformation-based motion modeling [91] is provided in the supplementary material.
@@
3D tracker / GT pose vs. Our object tracking module. In Tab. 3 (D), we use the 3D tracker [11, 66] to reconstruct dynamic objects and supervise motion learning, resulting in a 0.97 dB DPSNR decrease. Removing motion loss leads to a further 1.08 dB drop (Tab. 3 (E)), showing that our motion learning strategy compensates for tracking errors, though these errors still impact novel view synthesis, emphasizing the need for a more generalizable 2D tracker. Using ground-truth trajectories (Tab. 3 (F)) shows only a small gap, confirming the effectiveness of our 2D tracker in motion learning. See supplementary for tracking error analysis.
@@ 5 Conclusion
In this paper, we introduce a novel framework for robust dynamic 3D street scene reconstruction that eliminates the reliance on 3D object trackers. Addressing the generalization limitations of 3D trackers, we propose a robust object tracking strategy based on a 2D foundation model. Our framework also features a motion learning module within an implicit feature space to handle inevitable tracking errors by autonomously refining pose inaccuracies and recovering missed detections. Experiments on the Waymo-NOTR and KITTI datasets demonstrate its adaptability and superior performance. Limitations and failure cases are discussed in the supplementary material.
Acknowledgment
This work was supported by the National Key R&D Program of China under Grant 2018AAA0102801, PCL-CMCC Foundation for Science and Innovation Grant No. 2024ZY1C0040, and the National Research Foundation, Singapore, under its NRF-Investigatorship Programme (Award ID. NRF-NRFI09-0008).
Supplementary Material
6 Evaluation on Static Scenes in Waymo-NOTR Dataset
We evaluate our method on the static32 subset of the Waymo-NOTR dataset [92, 76], following the experimental setup of EmerNeRF [92] for novel view synthesis. As shown in Table 4, while our primary focus is on handling dynamic objects, our method also demonstrates robust performance in static scenes.
Method PSNR↑ SSIM↑ LPIPS↓
3DGS [37] 26.82 0.836 0.134
EmerNeRF [92] 28.89 0.814 0.212
S3Gaussian [32] 27.05 0.825 0.142
MARS [85] 27.63 0.848 0.193
Ours 28.72 0.857 0.092
Table 4: Comparison with state-of-the-art methods on the static32 subset of Waymo-NOTR dataset. StreetGS represents Street Gaussian [91]. The best and the second best results are denoted by pink and blue.
7 Runtime Analysis
As shown in Table 5, we evaluate the inference speed of our method and several state-of-the-art methods at a resolution of $960\times 640$ on the same device.
Method 3DGS [37] S3G [32] SG [91] Ours
Speed (FPS) 200 15 160 100
Table 5: Inference speed at $960\times 640$. S3G and SG represent S3Gaussian [32] and Street Gaussians [91] respectively.
8 Editing Examples
We provide editing demonstrations in Fig. 4. Gaussians corresponding to cars are associated at initialization and consistently maintained throughout the optimization process. This enables object editing by directly applying rigid transformations to the corresponding Gaussians.
Figure 4: Editing demonstrations on Waymo-NOTR.
9 Choice of 3D tracker
3D tracker PSNR↑ SSIM↑ LPIPS↓ DPSNR↑ DSSIM↑
Street Gaussians CasTrack [83, 82] 25.61 0.816 0.163 21.07 0.597
Street Gaussians VoxelNext [11] 26.98 0.838 0.149 24.62 0.742
Table 6: Ablation study on the choice of 3D tracker for Novel View Synthesis on the Waymo-NOTR dataset.
To further illustrate the generalization challenges of 3D trackers, we employ CasTrack [83, 82] as the 3D tracker for Street Gaussians [91], using the same detection and tracking algorithm as in the original paper. Since pretrained nuScenes weights are unavailable, we instead use weights pretrained on KITTI. As shown in Table 6, this change leads to a significant performance drop, indicating that merely changing the detection or tracking algorithm does not resolve generalization issues.
10 Tracking Errors Analysis
We evaluate 3D trajectories from 2D and 3D trackers on Waymo-NOTR, measuring translation (Euclidean) and rotation errors (clipped at $1m$ or $30^{\circ}$ ; with missing detections treated as max error). Our 2D tracker-based method is more robust. Error distributions are shown in Fig. 5; a qualitative comparison is shown in Fig. 1 in the main text.
Figure 5: Distributions of tracking errors.
11 Comparison with Rigid-transformation-based Motion Modeling
On Waymo-NOTR, we use 3D trajectories computed from 2D tracking to model vehicle motion via rigid transformations, following Street Gaussians (SG) [91] (see Tab. 7). While improved trajectories help, SG still underperforms—its joint optimization of pose and Gaussians struggles to converge in frames with large pose noise or missing detections, which in turn degrades the Gaussians across all frames. In contrast, we use HexPlane to enforce temporal smoothness by interpolating motion features from a spatio-temporal voxel grid. Since we remove trajectory guidance after the first 40% of training iterations, HexPlane can recover from noisy or missing detections by learning consistent motion from correctly tracked frames.
[…]
2D tracker + rigid transformation 27.69 0.850 25.04 0.758
3D tracker + rigid transformation 26.98 0.838 24.62 0.742
Ours (2D tracker + HexPlane) 28.85 0.867 25.58 0.779
Table 7: Comparison of rigid-transformation-based motion modeling [91] with our method on the Waymo-NOTR dataset.
12 Details of Loss Functions
As described in the main paper, the total loss function is expressed as:
$\begin{split}\mathcal{L}=&\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{tv}}\mathcal{L}_{\text{tv}}\\ &+\lambda_{\text{color-reg}}\mathcal{L}_{\text{color-reg}}+\lambda_{\text{motion}}\mathcal{L}_{\text{motion}}.\end{split}$ (9)
The components of the loss function are detailed below:
  1. 1.
    Photometric L1 Loss ($\mathcal{L}_{\text{rgb}}$): This L1 loss measures the photometric difference between the rendered image and the ground truth:
[…]
    where $\mathbf{I}_{\text{render}}$ and $\mathbf{I}_{\text{gt}}$ represent the rendered and ground truth images, respectively.
  2. 2.
    Structural Similarity Loss ($\mathcal{L}_{\text{ssim}}$): This loss evaluates the structural similarity between $\mathbf{I}_{\text{render}}$ and $\mathbf{I}_{\text{gt}}$:
    $\mathcal{L}_{\text{ssim}}=1.0-\text{SSIM}(\mathbf{I}_{\text{render}},\mathbf{I}_{\text{gt}}).$ (11)
  3. 3.
    Depth Loss ($\mathcal{L}_{\text{depth}}$): This L1 loss computes the difference between the rendered depth map $\mathbf{D}_{\text{render}}$ and the ground truth depth map derived from LiDAR data $\mathbf{D}_{\text{gt}}$:
[…]
    where $d=80$ is the predefined maximum depth used for normalization. Depth loss is calculated only for pixels with ground truth depth values between 0.01 and 80 meters.
  4. 4.
    Total Variation Loss ($\mathcal{L}_{\text{tv}}$): A grid-based total variation loss is employed to encourage smooth gradients for HexPlane feature grids, following K-Planes [70]:
[…]
    where $\operatorname{avg}$ denotes the average operator, $c$ is the plane index, and $i,j$ are indices on the plane resolution.
  5. 5.
    Color Regularization Loss ($\mathcal{L}_{\text{color-reg}}$): This L1 regularization loss minimizes the predicted color change $\Delta\mathcal{C}$ for each point to regularize the deformation network:
[…]
  6. 6.
    Motion Loss $\mathcal{L}_{\text{motion}}$: The motion loss is introduced in the main paper as,
    where $\mathcal{X}$ is the center position of a Gaussian in object $\mathcal{O}$.
The weights assigned to each loss component are: $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{ssim}}=0.1$, $\lambda_{\text{depth}}=1.0$, $\lambda_{\text{tv}}=0.1$, $\lambda_{\text{color-reg}}=0.01$, and $\lambda_{\text{motion}}=1.0$.
13 Limitation and Failure Cases
  • Our approach primarily focuses on modeling moving vehicles while using 4DGS [81] to model humans without explicit motion guidance. This leads to inaccurate human motion in some cases, which can be improved by incorporating human pose estimation as prior information (see Fig. 6). Incorporating human pose estimation techniques [88, 89] in future work could enhance human motion modeling.
  • Dynamic objects beyond humans and vehicles, such as animals and traffic lights, are currently treated as static objects. Further refinement is needed to more effectively distinguish between static and dynamic elements within a scene.
  • The proposed method requires per-scene optimization. A promising direction for future work is the development of a feed-forward approach for predicting generalizable 3D Gaussians.
Figure 6: Failure cases